OpenStudy (anonymous):

The length of a rectangle is 5 cm less than twice its width. The perimeter of the rectangle is 80 cm. What are the dimensions of the rectangle. Help?

OpenStudy (anonymous):

@AriPotta Do you think you are able to help me with this? :)

OpenStudy (aripotta):

P = 2L + 2W L = 2W - 5 substitute 2W - 5 for L in the perimeter formula: 80 = 2(2W - 5) + 2W

OpenStudy (aripotta):

solve for W

OpenStudy (anonymous):

So..80 = 4w - 10 + 2W?

OpenStudy (aripotta):

yes. now try to get W by itself

OpenStudy (anonymous):

Would I minus the 2W from the 4W?

OpenStudy (anonymous):

Oh wait nvm

OpenStudy (aripotta):

no, you would add them since they're both positive

OpenStudy (anonymous):

80 = -10 + 6W?

OpenStudy (aripotta):

yes, good job far :) a few more steps left

OpenStudy (aripotta):

good job so far*

OpenStudy (anonymous):

Umm..Would I take away 6 from both sides or add ten to both sides? I feel like I should take away 6.

OpenStudy (aripotta):

add 10

OpenStudy (anonymous):

so 80 = 16w?

OpenStudy (aripotta):

no you can't add 10 with 6w cuz they're not like terms. you wanna add 10 to both sides. you'd get 90 = 6w

OpenStudy (anonymous):

OhhH! 90 divided by 6..Which is 15..So 15 is the answer?

OpenStudy (aripotta):

yes :) W = 15 now to get L. L = 2W - 5. substitute 15 for W. L = 2(15) - 5

OpenStudy (anonymous):

25?

OpenStudy (aripotta):

yup. so your dimensions are 25 by 15

OpenStudy (aripotta):

great job :)

OpenStudy (anonymous):

YAY! :D Thank you

OpenStudy (aripotta):

no problemo